OpenStudy (anonymous):

Simplify the following expression. 3x − 4y + 3z − 3x − y + 2z

OpenStudy (anonymous):

3x+(-3x)=0 -4y+(-y)=-5y 3z+2z=5z So now you'll have -5y+5z

OpenStudy (anonymous):

yep